Asailboat travels 20 miles downstream in 3 hours. It returns in 4 hours. Find the speed of the sailboat in still water and the rate of the current. Plus how would you check your answer.
Answer by checkley77(12569)

D=RT
R=D/T
R=20/3=6.667 MPH. DOWN STREAM
R=20/4=5 MPH. UP STREAM
(6.667-5)/2=1.667/2=.833 MPH. FOR THE CURRENT.
6.667-.833=5.833 SPEED OF THE BOAT.
5+.833=5.833 DITTO.